### v1

# Kernel Search approach to solve the Minimum Spanning Tree Problem with conflicting edge pairs00footnotetext: Email addresses:

Abstract
The Minimum Spanning Tree Problem with Conflicts consists in finding the minimum conflict-free spanning tree of a graph, i.e., the spanning tree of minimum cost, including no pairs of edges that are in conflict. In this paper, we solve this problem using a tailored Kernel Search heuristic method, which consists in solving iteratively improved restrictions of the problem. The main novelty of the approach consists in using an independent set of the conflict graph within the algorithm. We test our approach on the benchmark instances and we compare our results with the ones obtained by other heuristics available in the literature.
Keywords: Minimum Spanning Tree, Conflicting Edges, Kernel Search, Cybersecurity, Disjunctive Constraints, Matheuristic
1 Introduction
[…]
Other heuristics are proposed in [14], where a Greedy Randomized Adaptive Search Procedure (GRASP) method coupled with adaptive memory programming is presented, and in [5], where a Multiethnic Genetic Algorithm (MEGA) is designed for the Minimum Conflict Weighted Spanning Tree variant, where the primary objective is to find a spanning tree with the minimum conflicting edge-pairs, and the second one is to minimize the weight of the tree, given it is conflict-free. Comparative studies demonstrated MEGA superiority over the tabu search algorithm proposed in [35]. In [9], another genetic algorithm is proposed, which is computationally better than both the tabu search and MEGA. Recently, in [11] an approach consisting of two phases of metaheuristic techniques is proposed for the same variant of MSTC: first, a hybrid artificial bee colony algorithm is used to detect a conflict-free spanning tree, and, second, a iterated local search that starts with this conflict-free solution is used to find the MSTC of minimum weight.
In [32], a general preprocessing method and a branch and cut algorithm are proposed with the generation of cutting planes corresponding to subtour elimination constraints and odd-cycle inequalities. Another branch and cut method for the MSTC is proposed in [6], where a new set of valid inequalities for the problem is introduced. To obtain upper bounds that help prune the search tree, the authors of [6] use the genetic algorithm proposed in [5]. As in [35], in [7], the Lagrangian relaxation of the MSTC is studied and solved with an ad hoc dual ascent procedure. In order to obtain lower bounds for MSTC, the authors of [8] proposed a subgradient method that uses a new rule to calculate the step size based on retaining some information from the previous iteration.
In this paper, we propose a new solution approach to the MSTC based on the so-called Kernel Search (KS) method [1]. It falls within the category of general-purpose methods for solving Mixed-Integer Linear Programming (MILP) problems. Initially employed in addressing challenges like the multidimensional knapsack problem [1] and a portfolio selection problem [2], it has been used also to solve, among others, facility location problems with or without conflicts [25, 19, 21], as well as the multivehicle inventory routing problem [3]. The main purpose of this algorithm is to obtain a feasible solution, hopefully of good quality, from a small set of promising variables of the MILP. This subset of variables constitutes the so-called kernel set that is initially built using the information provided by solving the Linear Programming (LP) relaxation of the original MILP. Throughout the algorithm, new promising variables to add to the kernel set may be found by solving MILP subproblems of increasing size.
An important novelty in our approach relies on the use of independent sets of proper subgraphs of the conflict graph, within the KS algorithm. This idea is derived from the observation that the set of edges of any feasible solution of MSTC corresponds to an independent set in the conflict graph, and we believe that it can be applied to other optimization problems with conflicts, whenever the conflict graph is used.
The rest of the paper is organized as follows. In Section 2, we present the Integer Linear Programming (ILP) formulation of MSTC we consider. In Section 3, the tailored kernel search approach we use to solve this formulation is proposed and discussed. In Section 4, we test the KS approach on the benchmark instances and compare our results with the ones obtained by other heuristics available in the literature. Finally, Section 5 concludes the paper.
2 ILP Formulation of MSTC and notation
Let us consider an undirected graph $\mathcal{G}=(V,E)$, where $V=\{1,\dots,n\}$ is the set of nodes and $E=\{e_{1},\dots,e_{m}\}$ the set of edges. The edge $e_{i}$ can be defined also by its two endpoints $u,v\in V$. For a given subset of nodes $S\subseteq V$, $E(S)=\{\{u,v\}\in E~{}|~{}u,v\in S\}$ is the set of edges having both the two endpoints in $S$. Let $C$ be the set of conflicting edge pairs, which we call conflict set: $\;C=\{(e_{i},e_{j})~{}:~{}e_{i},e_{j}\in E,\text{ and }e_{i}\text{ is in conflict with }e_{j}\}.$
The general formulation for MSTC uses the following binary variables, defined for each $i$ s.t. $e_{i}\in E:$
$x_{i}=\begin{cases}1&\text{if edge $e_{i}$ is included in the tree}\\ 0&\text{otherwise.}\end{cases}$
The ILP formulation of the MSTC, coming from the Subtour Elimination formulation of the classical MST, reads:
$\min\limits_{x}$ $\sum_{i:e_{i}\in E}w_{i}x_{i}$ (1a)
s.t. $\;\sum\limits_{i:e_{i}\in E}x_{i}=n-1$ (1b)
$\sum\limits_{i:e_{i}\in E(S)}x_{i}\leq|S|-1$ $\forall\,S\subset V,|S|\geq 3$ (1c)
$x_{i}+x_{j}\leq 1$ $\forall\,i,j:(e_{i},e_{j})\in C$ (1d)
$x\in\{0,1\}^{|E|}$ (1e)
The objective function (1a) represents the cost of the spanning tree, which we want to minimize. Constraint (1b) imposes that $n-1$ edges must be selected since a spanning tree must include exactly $|V|-1$ edges. The exponentially many constraints (1c) are needed to ensure that no subtour is contained in the selected edges, i.e., no subgraph induced by $S$ contains a cycle. Finally, constraints (1d) impose that, for each pair of edges in conflict, at most one edge of such pair can be selected.
In the following, we denote by LP-MSTC the LP relaxation of model (2), obtained by relaxing constraints (1e). Given a subset of edges $F\subseteq E$, we further denote by MSTC$(F)$ the ILP restriction of problem (2) obtained by restricting the set of variables to the edges in $F$, i.e., adding constraints
$x_{i}=0\quad\forall i\in E\setminus F$
to formulation (2). As before, LP-MSTC$(F)$ denotes the LP relaxation of MSTC$(F)$.
Furthermore, to indicate the $k$-th element of a set $A$, we use the notation $A[k]$, with $A[1]$ representing the first element of $A$, and to refer to the first $k$ elements of set $A$, we use the notation $A[1:k]$. Instead, given a graph $\mathcal{H}$, a subset of its nodes $\tilde{V}$, and a subset of its edges $\tilde{E}$, we denote by $\mathcal{H}[\tilde{V}]$ the induced subgraph of $\mathcal{H}$ whose node set is $\tilde{V}$ and whose edge set consists of all of the edges of $\mathcal{H}$ that have both endpoints in $\tilde{V}$. Whereas, we denote by $\mathcal{H}[\tilde{E}]$ the induced subgraph of $\mathcal{H}$ whose edge set is $\tilde{E}$ and whose node set consists of all of the endpoints of $\tilde{E}$ in $\mathcal{H}$.
To conclude this section, we recall some important definitions. The set $N(u)$ of the neighbors of node $u\in V$ is the set of all the nodes $i\in V\setminus\{u\}$ such that it exists an edge $\{u,v\}\in E$. An independent set is a set of nodes in a graph, no two of which are adjacent. Given the graph $\mathcal{G}=(V,E)$, its conflict graph $\hat{\mathcal{G}}=(\hat{V},\hat{E})$ is a graph which has a node for each edge in $\mathcal{G}$, and where we represent each conflict constraint by an edge connecting the corresponding vertices in $\hat{\mathcal{G}}$. Note that each conflict-free spanning tree in $\mathcal{G}$ is a subset of $\hat{V}$ which corresponds to an independent set of $\hat{\mathcal{G}}$.
3 Kernel Search approach
In this section, we describe the KS framework we use to solve MSTC. Given the mathematical formulation (2) of MSTC, the starting point of the KS algorithm consists in identifying a subset of promising variables of this model, i.e., variables that have a high probability of being non-zero in the optimal solution of the problem. These variables are used to form a subset called kernel set, while the remaining ones are partitioned into several subsets called buckets. We define $\Lambda^{0}$ as the set of indices of the variables in the initial kernel set, and $B_{i}$ as the set of indices of the variables in the $i$-th bucket, for $i=1,\dots,b$.
The partitioning of the indices of the variables into the kernel set and the buckets is carried out by KS according to the optimal solution of LP-MSTC, an MST solution (not necessarily conflict-free) returned by the GRASP algorithm proposed in [14], as well as an independent set of a proper subgraph of the conflict graph $\hat{\mathcal{G}}$. After the construction of the kernel set and the buckets, KS seeks an initial feasible solution by solving the MSTC restricted to the variables having indices inside the kernel set only, i.e. solving MSTC($\Lambda^{0}$). For this reason, the initial kernel size $|\Lambda^{0}|$ should be, at the same time, large enough to contain as many variables as possible among the ones that are likely to be part of an optimal solution, and not too large to ensure that MSTC($\Lambda^{0}$) can be efficiently solved to optimality.
The kernel set is then updated throughout the KS algorithm ($\Lambda^{k}$ is the kernel set at the end of iteration $k$), by taking into account the buckets. At each iteration, indeed, the MILP formulation, restricted to the current kernel set and a bucket from the sequence, is solved, trying to obtain solutions of increasing quality and thus new variables to insert into the kernel set.
To sum up, the KS algorithm is divided into two phases:
  • Initialization phase: The initial set $\Lambda^{0}$ and the bucket sequence =$\mathcal{B}=\{B_{1},\dots B_{b}\}$ are constructed. Then, the MSTC($\Lambda^{0}$) is solved and, if a feasible solution is found, its objective function value is set as the current upper bound $w^{*}$ of the optimal solution we are looking for.
  • Extension phase: At each iteration $k$, new variables are added to the kernel set iterating (eventually more than once) over the buckets, and the MSTC restricted on the so-obtained enlarged kernel is solved. If this restriction is feasible, its optimal value is denoted by $w^{k},$ and eventually used to update the value of $w^{*}$, and its optimal solution is denoted by $x^{k},$ and eventually used to update the kernel set and the buckets.
Let us describe in more detail how these two phases work.
3.1 Initialization phase
The construction of starting kernel set $\Lambda^{0}$ and of set $\mathcal{B}$ is carried out by Algorithm 2. The algorithm takes as input two positive parameters $K$ and $d$. First of all, the optimal solution of LP-MSTC is computed (line 1). The variables that have a positive value in the solution of LP-MSTC are then sorted according to their values, in non-increasing order (line 2). Being $N$ the set of edge indices associated with the positive variables ordered in this way, we set $K$ in such a way that it is not greater than the cardinality of $N$ (line 3). We define the subset of the first $K$ elements of $N$ as $N_{K}$ and the set of the pairs of edges in $N_{K}$ which are in conflict as $C_{K}$, i.e., $C_{K}=\{(i,j)~{}:~{}i,j\in N_{K},\text{ and }(e_{i},e_{j})\in C\}$ (line 4). The graph $\hat{\mathcal{G}}_{K}=(N_{K},C_{K})$ (subgraph of the conflict graph $\hat{\mathcal{G}}$) is taken into account. In line 5, a feasible solution $I$ of the maximum independent set problem on this graph $\hat{\mathcal{G}}_{K}$, with the guarantee of not generating cycles in the starting graph $\mathcal{G}$, is computed through a greedy procedure presented in Algorithm 1, by setting $W=N_{K}$ and $\mathcal{H}=\hat{\mathcal{G}}_{K}$.
We use the maximum independent set problem to find the most promising variables by reason of the fact that a spanning tree $(V,E^{\prime})$ is a feasible solution for the MSTC problem iff $E^{\prime}$ is an independent set of the conflict graph. Notice that this is a necessary but not sufficient condition to obtain a feasible solution for the MSTC. Since Algorithm 1 selects edges that are not in conflict with each other and that do not generate cycles in the starting graph $\mathcal{G}$, then these edges induce either a feasible solution of the MSTC or a set of conflict-free subtrees of $\mathcal{G}$.
This means that, by selecting the variables according to the solution of Algorithm 1, we increase the probability of including edges in the kernel set that can generate feasible solutions for the MSTC. It is worth noting that, actually, the idea of using the maximum independent set can be extended to any optimization problem with conflicts because its feasible solutions must be independent sets of the conflict graph.
Input: Graph $\mathcal{H}$, and node set $W$. Output: Independent set $I$ of $\mathcal{H}$. 1 Set $I=\emptyset,\bar{V}=\emptyset,\bar{E}=\emptyset$. 2 while $W\neq\emptyset$ do 3       Select a node $i\in W$ with minimum degree in $\mathcal{H}[W]$, and let $e_{i}=\{u,v\}\in E$ be the edge of $\mathcal{G}$ associated to $i$. 4       Set $\bar{V}=\bar{V}\cup\{u,v\}$ and $\bar{E}=\bar{E}\cup\{e_{i}\}.$ 5       if  $\mathcal{G}[\bar{E}]$ does not contain cycles then 6            Set $I=I\cup\{i\}.$ 7             Being $W_{i}$ the set of neighbors of $i$ in $\mathcal{H}$, including $i$, set $W=W\setminus W_{i}.$ 8      else 9             Set $W=W\setminus\{i\}.$ 10       end if 11       12 end while return $I$ Algorithm 1 Algorithm to compute an independent set of the graph $\mathcal{H}$
Once the set $I$ is computed through Algorithm 1, the GRASP algorithm proposed in [14] is executed (line 5 as well). This multistart metaheuristic consists of two phases: (i) a constructive phase, which creates an initial solution using the adaptive memory (i.e., taking into account a measure of how frequently an edge occurs in the best solutions generated in previous iterations); (2) a local search procedure, moving from the current solution to neighbor ones, trying to find better solutions. The output of this algorithm is a tree $(V,\bar{E})$, not necessarily without conflicts.
Given the two edge sets $I$ and $\bar{E}$, coming from Algorithm 1 and GRASP method, respectively, in line 6 the initial kernel set $\Lambda^{0}$ is populated with the indices of the variables associated with the edges in $I\cup\bar{E}$, and the remaining indices are used to define set $L$. In order to populate the buckets, the elements of set $L$ are sorted in non-increasing order of the number of conflicts with the edges in $\Lambda^{0}$ in line 7. The variables corresponding to edges with no conflict with the edges in $\Lambda^{0}$ are sorted in the following way: first, the positive variables in the solution of the LP-MSTC (with indices in $L\cap N$) in non-increasing order of their values, second the remaining ones (with indices in $L\cap\{E\setminus N\}$) in non-increasing order of their reduced cost coefficients. Then, in line 8, these variables, ordered as explained, are equally distributed to populate the buckets $B_{1},\dots,B_{b},$ each of them of cardinality $d$ (except for the last bucket, the size of which could be smaller). The set $\mathcal{B}$ is defined as the set of buckets. In line 9, the initialization phase is concluded by solving the MSTC($\Lambda^{0}$) formulation, if feasible. The algorithm returns its optimal value $w^{0}$ and its optimal solution $x^{0}$, as well as sets $\Lambda^{0}$ and $\mathcal{B}$.
Input: Graph $\mathcal{G}=(V,E),$ two scalars $K>0,d>0$. Output: Initial kernel set $\Lambda^{0}$, initial bucket set $\mathcal{B}$, the first MSTC solution found $x^{0}$, and its value $w^{0}.$ 1 Compute the optimal solution $\bar{x}$ of the LP-MSTC model. Define set $N:=\{i\in E:\bar{x}_{i}>0\}$. 2 Sort indices $i\in N$ in non-increasing order of $\bar{x}_{i}$ values. 3 Set $K=\min\{K,|N|\}$. 4 Define $N_{K}:=N[1:K]$, and $C_{K}:=\{(i,j)~{}:~{}i,j\in N_{K},\text{ and }(e_{i},e_{j})\in C\}$. 5 Compute the set $I$ through Algorithm 1, and the set $\bar{E}$ through the GRASP algorithm [14]. 6 Define $\Lambda^{0}=I\cup\bar{E}$, and $L:=E\setminus\Lambda^{0}$. 7 Sort indices $i\in L$ in non-increasing order of number of conflicts with the edges in $\Lambda^{0}$. 8 Populate buckets $B_{i}$ by taking $d$ variables at a time from $L.$ Being $b$ the number of buckets, define =$\mathcal{B}=\{B_{1},\dots,B_{b}\}.$ 9 if MSTC$(\Lambda^{0})$ is feasible then 10       Compute the optimal solution $x^{0}$ of MSTC$(\Lambda^{0})$, and its optimal value $w^{0}$. 11else 12      $x^{0}=0$, and $w^{0}=+\infty$ 13 end if return $\Lambda^{0},\mathcal{B},x^{0},w^{0}.$ Algorithm 2 KS algorithm: Initialization Phase
3.2 Extension phase
With the initial kernel set $\Lambda^{0}$, the initial set of buckets $\mathcal{B}$, as well as the initial solution $x^{0}$ (which is the initial best-known solution $x^{*}$), if any, and its value $w^{0}$ (which is the best known upper bound $w^{*}$), returned by the initialization phase (Algorithm 2), the extension phase, described in Algorithm 3, starts. We recall that, if the initialization phase does not provide any feasible solution, then $x^{0}$ is set to 00 and $w^{0}$ is set to +$\infty$+ ∞ (see Step 12 of Algorithm 2).
During this phase, both the kernel set and the set of buckets are updated. The while loop of line 2 manages the number of runs that the algorithm carries out on the buckets and, according to the value of the index $p$, how the buckets are combined to create the restricted problems (lines 3-4). In particular, when $p=1$, the restricted problems are generated by considering one bucket from the original set of buckets $\mathcal{B}$ at a time; if $p>1$, multiple buckets, properly selected, are considered at a time. The list of enlarged buckets $\tilde{\mathcal{B}}$ is updated, according to the $p$ value, by Algorithm 4 described in the following.
At each iteration $k$ of the while loop at line 7, following the Kernel Search framework, the kernel set is, eventually, updated by embedding in it variables coming from a bucket. In particular, at iteration $k$, the algorithm solves the restricted problem $MSTC(\Lambda^{k-1}\cup\tilde{B}_{k})$ (line 9) whose variables are the ones inside the current kernel set $\Lambda^{k-1}$ and the current bucket $\tilde{B}_{k}$ (coming from Algorithm 4). This restricted problem is solved with the additional constraints $\sum_{i\in\tilde{B}_{k}}{x_{i}}\geq 1$ and $\sum_{i:e_{i}\in E}w_{i}x_{i}\leq w^{*}$ to ensure, respectively, that at least one new variable, from the current bucket, is used, and that the value of the new solution $x^{k}$ is at least as good as the one of the current incumbent $x^{*}$.
If $MSTC(\Lambda^{k-1}\cup\tilde{B}_{k})$ is feasible (i.e., its optimal value $w^{k}$ is better than or equal to the current upper bound $w^{*}$), then the incumbent solution $x^{*}$ and its value $w^{*}$ are updated accordingly (line 10). The subset $\tilde{B}_{k}^{+}$ of variables of the current bucket $\tilde{B}_{k}$ which are nonzero in the solution $x^{k}$ is added to the kernel set (line 12) and removed from the corresponding bucket (line 13), i.e., the indices belonging to subset $\tilde{B}_{k}^{+}$ are added to $\Lambda^{k-1}$, and removed from $\tilde{B}_{k}$. Finally, the parameter $\bar{k}$, which is needed for the stopping criterion of the algorithm, is set to zero, because a new incumbent solution was found (line 14). If, instead, the restriction is infeasible, i.e., $w^{k}\geq w^{*}$, we increase parameter $\bar{k}$ (line 16). Indeed, if $\bar{k}\geq\delta,$ i.e., for at least $\delta$ steps, the solution has not been improved, and at least a feasible solution has been found so far (line 19), the algorithm updates the parameters $k$ and $p$ (lines 20–21) to terminate the two while loops at lines 2 and 7.
In the last step, the algorithm returns the best found solution $x^{*}$ and its objective function value $w^{*}$ (line 27).
Input: The solution $x^{0}$, its value $w^{0},$ sets $\Lambda^{0},\mathcal{B}=\{B_{1},\dots,B_{b}\},$ and scalars $\delta>0,P>0.$ Output: Best solution found $x^{*}$ and its value $w^{*}$. 1 Let ~$\tilde{\mathcal{B}}=\mathcal{B},x^{*}=x^{0},w^{*}=w^{0}$, $p=1$, $\bar{k}=0$. 2 while $p\leq P$ do 3       if $p>1$ then 4            $\tilde{\mathcal{B}}=$ Algorithm 4$(\tilde{B},\Lambda^{0}).\quad$ //Update the set of enlarged buckets 5       end if 6      Set $k=1$. 7       while $k\leq b$ do 8             if MSTC$(\Lambda^{k-1}\cup\tilde{B}_{k})$ is feasible then 9                  Compute the optimal solution $x^{k}$ of MSTC$(\Lambda^{k-1}\cup\tilde{B}_{k})$, and its optimal value $w^{k}$. 10                   Set $x^{*}=x^{k}$, and $w^{*}=w^{k}$. 11                   Set $\tilde{B}_{k}^{+}=\{i\in\tilde{B}_{k}:x^{k}_{i}>0\}$. 12                   Set $\Lambda^{k}=\Lambda^{k-1}\cup\tilde{B}_{k}^{+}$. 13                   Set $\tilde{B}_{k}=\tilde{B}_{k}\setminus\tilde{B}_{k}^{+}$. 14                   Set $\bar{k}=0$. 15            else 16                  Set $\bar{k}=\bar{k}+1$. 17             end if 18            Set $k=k+1$. 19             if $\bar{k}\geq\delta$ and $w^{*}<+\infty$ then 20                   Set $k=b+1$. 21                   Set $p=P+1$. 22             end if 23             24       end while 25      Set $p=p+1$. 26       Set $\Lambda^{0}=\Lambda^{k}$. 27 end while 28return $w^{*},x^{*}$ Algorithm 3 KS algorithm: Extension Phase
Let us now describe the Algorithm 4 which is called at the beginning of each iteration of the while loop, at line 2, of Algorithm 3. It starts by initializing the set $\hat{\mathcal{B}}$, used to define the new bucket set, to the empty set (line 1). The following two for loops (lines 2–8) are used to select all the possible couples of buckets from $\tilde{\mathcal{B}}$. For each couple $A_{lt}$, the algorithm builds the conflict subgraph $\mathcal{H}=(Y,C_{Y})$, the nodes of which are all the edges in $\Lambda^{0}\cup A_{lt}$ (line 5). Two nodes $e_{i}$ and $e_{j}$ of $Y$ are connected if their corresponding edges in $\mathcal{G}$ are in conflict, i.e. $(e_{i},e_{j})\in C$. At line 6, an independent set of $\mathcal{H}$ is computed by using Algorithm 1. The cardinality of this independent set gives us an estimation of the compatibility between the edges inside the kernel set and the ones inside $A_{lt}$. Indeed, the greater this cardinality is, the higher it will be the probability of producing feasible solutions of MSTC by solving the restricted problem $MSTC(\Lambda^{0}\cup A_{lt})$. For this reason, given the set $\hat{A}$ of all the bucket pairs $A_{lt}$, these pairs are sorted, in non-increasing order of the cardinality of the associated sets $I_{lt}$ (lines 9-10). At each iteration $k$ of the while loop at line 12 (with $k$ initialized to $1$ in line 11), the $k$-th couple of $\hat{A}$, say $A_{lt}=\tilde{B_{l}}\cup\tilde{B_{t}}$, is considered. The buckets $\tilde{B_{l}}$ and $\tilde{B_{t}}$ are removed from the set $\tilde{\mathcal{B}}$ and all the couples containing either $\tilde{B_{l}}$ or $\tilde{B_{t}}$ are removed from $\hat{A}$ (line 13). Furthermore, the new enlarged bucket $\tilde{B_{l}}\cup\tilde{B_{t}}$ is added to $\hat{\mathcal{B}}$ (line 14). If the number of current buckets is odd, the last bucket is added to $\hat{\mathcal{B}}$ (line 18), i.e., it will be again considered alone in Algorithm 3. In the last two lines, the algorithm assigns $\hat{\mathcal{B}}$ to $\tilde{\mathcal{B}}$ and returns this last set of enlarged buckets.
Input: Sets ~=$\tilde{\mathcal{B}}=\{\tilde{B}_{1},\dots,\tilde{B}_{b}\}$, $\Lambda^{0}$. Output: Updated set of enlarged buckets ~=$\tilde{\mathcal{B}}=\{\tilde{B}_{1},\dots,\tilde{B}_{b}\}$, with $b=|\tilde{\mathcal{B}}|$. 1 Let $\hat{\mathcal{B}}=\emptyset$. 2 for $l\in\{1\leq b-1\}$ do 3       for $t\in\{l+1\leq b\}$ do 4             Define set $A_{lt}=\tilde{B}_{l}\cup\tilde{B}_{t}$. 5             Define the sets $Y=\Lambda^{0}\cup A_{lt}$, and $C_{Y}=\{(i,j)~{}:~{}i,j\in Y,\text{ and }(e_{i},e_{j})\in C\}$. 6             Compute the independent set $I_{lt}$ giving in input to Algorithm 1 set $W=Y$ and graph $\mathcal{H}=(Y,C_{Y})$. 7             8       end for 9       10 end for 11Define the set $\hat{A}$ as the set of all the obtained bucket pairs $A_{lt}$. 12 Order set $\hat{A}$ elements in non-increasing order of cardinality of the associated sets $I_{lt}$. 13 Set $k=1$. 14 while $\hat{A}\neq\emptyset$ do 15       Being $\hat{A}[k]=A_{lt}$, set $\hat{A}=\hat{A}\setminus\{A_{ij}\}\,\forall\,i,j~{}:~{}i=l\vee j=t$, and ~=~$\tilde{\mathcal{B}}=\tilde{\mathcal{B}}\setminus\{\tilde{B}_{l}\cup\tilde{B}_{t}\}$. 16       Set ^=^$\hat{\mathcal{B}}=\hat{\mathcal{B}}\cup\hat{A}[k]$. 17       Set $k=k+1$. 18       19 end while 20if $-~\tilde{\mathcal{B}}\neq\emptyset$ then 21      Set ^=^~$\hat{\mathcal{B}}=\hat{\mathcal{B}}\cup\tilde{\mathcal{B}}[1]$. 22       23 end if 24Set $\tilde{\mathcal{B}}=\hat{\mathcal{B}}.$ return $-~\tilde{\mathcal{B}}$ Algorithm 4 Population of the enlarged bucket set $\tilde{\mathcal{B}}$.
3.3 Valid inequalities
In this section, we describe some valid inequalities for the MSTC, taken from [6], that we add to our model in order to speed up the resolution of the restricted problems solved throughout the kernel search algorithm.
The first type of inequalities, defined degree-cut inequalities, explicitly impose that, for each node of the graph, at least one incident edge should be selected in the solution:
$\sum\limits_{i:e_{i}\in N(u)}x_{i}\geq 1\quad\forall u\in V.\vspace*{-2mm}$ (2)
These constraints are not really useful when dealing with integer solutions. However, when solving an LP relaxation of the original formulation, these may help in finding a better bound of the optimal value. These $n$ inequalities are added to the MSTC model (2) as a-priori constraints.
The conflict-cycle inequalities are a stronger version of the subtour elimination constraints (1c) obtained by taking into account also the conflicts among the edges. Indeed, given a cycle $P\subset E$, constraints (1c) impose that the edges belonging to $P$ selected in any feasible solution of the problem should be at most $|P|-1$ (not all of them can be selected, otherwise a subtour is included in the solution). If an edge $e_{c}$ outside the cycle $P$ is in conflict with two edges belonging to the cycle $P,$ then in any feasible solution of MSTC, the number of selected edges in $P\cup\{e_{c}\}$ should be still less than $|P|-1$: no cycle, as well as no pairs of edges in conflict, should be included in the solution. This is imposed by the following inequalities:
$\sum\limits_{i:e_{i}\in P}x_{i}+x_{c}\leq|P|-1\quad\forall\text{ cycle }P\subset E,e_{c}\in E\setminus P.\vspace*{-2mm}$ (3)
These inequalities, as well as the classical subtour elimination constraints (1c), are exponentially many, thus a separation procedure should be implemented for them. In particular, the separation problem for constraints (1c), which consists in finding the violated constraints of type (1c) by a given solution, corresponds to solving a maximum-flow (minimum-cut) problem in a correspondingly defined capacitated graph as described in [28]. As regards conflict-cycle inequalities (3), the heuristic separation procedure proposed in [6] is implemented.
4 Computational Tests
In this section, we report the results obtained by solving the instances proposed in [35] [^1] Instances available at http://ic.ufal.br/professor/rian/Instances/Instances-MCFST-Zhang.zip., as well as the ones proposed in [6] [^2] Instances available at http://www.dipmat2.unisa.it/people/carrabs/www/DataSet/CMST.zip., with our Kernel Search Method (Algorithms 23). In the following, we refer to the instances proposed in [35] as ZKP instances and to the ones proposed in [6] as CCPR instances. The ZKP instances are generated by using different values for nodes, edges, and number of conflicts and they are classified into two types. By construction, there exists at least one conflict-free solution for all Type 2 instances, while this is not the case for Type 1 instances. In the computational tests, we consider only instances where at least the existence of a feasible solution is guaranteed. The CCPR instances have a number of nodes in $\{25,50,75,100\}$ and a number of edges such that the density of the graph $m(m-1)/2$ is equal to $0.2$, $0.3$, or $0.4$. The number of conflicting pairs $|C|$ is equal to 1%, 4%, 7% of the graph density. For each combination of these parameters (scenario), the authors generated 5 instances by varying the seed used for the generation of random numbers, obtaining a total of 180 instances.
We conducted an extensive parameter tuning process for the algorithm, exploring a range of values for key parameters, and selecting the ones yielding better results. In particular, we set a time limit of 3600 seconds for the whole algorithm and a time limit for each subproblem of 350 seconds for ZKP instances and 420 seconds for CCPR instances. The computing time may exceed the time limit of at most 350 seconds for the ZKP and 420 seconds for the CCPR instances, if the solution of the last subproblem starts after 3250 or 3180 seconds, respectively.
Furthermore, in our tests, we set the starting values of $K,$ and $d$:
$K=\lfloor 1.2(|V|-1)\rceil\quad d=\lfloor 0.1(|E|-K)\rceil,$
which are then updated during the initialization phase, in such a way that the starting kernel set contains a sufficient number of variables to eventually provide a feasible solution to the problem (i.e., $n-1$) plus possible alternatives, and the subproblems solved at each iteration over the kernel set and the buckets are not excessively large. Instead, the maximum number of outer iterations $P$ is set to $1$, for the ZKP instances, and to $4$, for the CCPR instances. Furthermore, we set the parameter $\delta$, indicating the maximum number of iterations the KS can perform without improving, to $\lfloor 0.3\cdot b\rfloor$ (it changes at every iteration $p$, since $b$ changes along these iterations).
Tables 1-2 reports the results obtained by testing the KS method on Type 1 and Type 2 ZKP instances (respectively) proposed in [35] (for which at least a feasible solution is found), compared to the ones obtained by using the following heuristic approaches:
  • the Multi-Ethnic Genetic Approach (MEGA) proposed in [5];
  • the Lagrangian-based heuristic HDA+ proposed in [7], which provides a lower bound on the optimal value of the problem by solving the Lagrangian dual of the MSTC, and an upper bound through a local search procedure;
  • the metaheuristic introduced in [14], GRASP-AM which we also use within our Kernel Search initialization phase;
  • the Two Phases Metaheuristic Technique (TPMT) proposed in [11], which combines a hybrid artificial bee colony algorithm, and an iterated local search;
  • the classic Kernel Search algorithm (ClassicKS) as introduced in [1] – never used to solve the MSTC problem – which uses neither the independent set nor the GRASP. It performs at most $P=4$ iterations over the buckets for both types of instances, merging them two by two at each iteration. The initial kernel set size is $K$, and the initial bucket size is $d$. The stopping criterion is the same used in our KS algorithm, with $\delta=\lfloor 0.3\cdot b\rfloor$.
Each subproblem involved in the KS and ClassicKS algorithms is solved using the state-of-the-art solver Cplex. They are coded in Python on an OSX platform, running on an Intel(R) Core(TM) i7-10700K CPU 3.80GHz with 32 GB of RAM, equipped with the IBM ILOG CPLEX (Version identifier: 22.1.0.0) solver (multi-thread mode). As concerns the time, in order to have a fair comparison, the CPU time reported in the literature for each method is properly scaled according to the Whetstone benchmark [33].
Instance MEGA HDA+ GRASP-AM TPMT ClassicKS KS
ID $n$ $m$ $|C|$ UB_best time gap time gap time gap time gap time gap time gap
1 50 200 199 $*{}^{*}$ 0.46 0.00 0.22 2.54 0.13 0.00 0.63 0.00 0.57 0.00 0.54 0.00
2 50 200 398 $*{}^{*}$ 0.44 0.00 0.29 1.30 0.13 0.00 0.79 0.00 16.42 0.00 0.85 0.00
3 50 200 597 $*{}^{*}$ 0.40 0.00 0.36 28.24 0.13 0.00 1.12 0.00 9.33 3.82 0.84 0.00
4 50 200 995 $*{}^{*}$ 0.42 0.91 0.51 19.86 0.13 0.00 1.68 0.00 1.49 30.06 0.67 0.00
5 100 300 448 $*{}^{*}$ 1.54 1.16 0.76 1.44 1.34 0.00 4.80 0.00 135.72 0.00 15.04 0.00
6 100 300 897 $*{}^{*}$ 1.16 7.72 1.38 100 1.21 0.00 5.27 0.07 377.76 0.67 4.72 0.00
7 100 500 1247 $*{}^{*}$ 3.33 0.00 1.90 0.61 1.30 0.00 4.68 0.00 207.63 0.16 9.13 0.00
8 100 500 2495 $*{}^{*}$ 3.29 3.37 3.53 36.97 3.22 0.00 10.81 0.00 1016.28 1.62 10.90 0.00
9 100 500 3741 7523 2.39 1.89 4.38 100 3.06 0.00 15.79 0.00 3944.75 100 5.18 0.00
13 200 600 1797 13831 46.58 8.60 5.51 100 23.18 1.67 42.35 0.00 3881.71 100 38.02 -0.06
16 300 800 3196 21480 67.62 2.75 12.17 100 38.68 0.89 29.36 0.00 1591.68 8.20 54.47 -0.18
Average values 11.60 2.40 2.82 44.63 6.59 0.23 10.66 0.01 1020.30 22.23 12.76 -0.02
%best 36.4 0 81.8 72.7 27.3 100
Table 1: Comparison of the results on feasible Type 1 ZKP instances proposed in [35].
[…]
The results of the average values line show that all the algorithms but ClassicKS are very fast because, on average, most of them require less than 13 seconds. The fastest algorithm is HDA+, which requires only 2.82 seconds but is the less effective one with an average gap of 44.63%. On the contrary, the most effective algorithm is KS which always finds the optimal solution, in the instances where it is known, and, for the last two instances, it reports a solution better than the best one known in the literature. Its final average gap is equal to -0.02% and these results are obtained, on average, in 12.76 seconds with a peak of 54 seconds. The effectiveness of KS in these instances is further certified by %best line where its value is equal to 100% since it always finds a solution that is at least as good as the best one known in the literature. Notice that it is the only algorithm to obtain this result. Similar results are obtained by TPMT and GRASP-AM which are faster than KS but show a higher average gap. Moreover, their %best value is equal to 72.7% and 81.8%, respectively. It is worth noting the huge difference between the results of ClassicKS and KS algorithms. ClassicKS is the slowest algorithm of the set and only in three cases, it shows a gap equal to 0%, while its average gap is very high (22.23%). These preliminary results already show the impact that the use of the independent set strategy and the GRASP algorithm have on the effectiveness and performance of ClassicKS.
Instance MEGA TPMT ClassicKS KS
ID $n$ $m$ $|C|$ UB_best time gap time gap time gap time gap
24 50 200 3903 $*{}^{*}$ 0.30 0.00 0.04 0.00 0.16 0.00 0.16 0.00
25 50 200 4877 $*{}^{*}$ 0.30 0.00 0.06 0.00 0.18 0.00 0.19 0.00
26 50 200 5864 $*{}^{*}$ 0.33 0.00 0.06 0.00 0.37 0.00 0.38 0.00
27 100 300 8609 $*{}^{*}$ 1.21 0.00 0.12 0.00 0.01 0.00 0.00 0.00
28 100 300 10686 $*{}^{*}$ 1.08 0.00 0.16 0.00 0.00 0.00 0.00 0.00
29 100 300 12761 $*{}^{*}$ 1.11 0.00 0.18 0.00 0.00 0.00 0.00 0.00
30 100 500 24740 $*{}^{*}$ 2.12 0.00 0.63 0.00 36.49 0.00 39.33 0.00
31 100 500 30886 $*{}^{*}$ 2.24 0.00 0.46 0.00 29.72 0.00 23.89 0.00
32 100 500 36827 $*{}^{*}$ 2.26 0.00 0.45 0.00 34.78 0.00 36.23 0.00
33 200 400 13660 $*{}^{*}$ 4.66 0.00 0.36 0.00 0.00 0.00 0.00 0.00
34 200 400 17089 $*{}^{*}$ 4.81 0.00 0.37 0.00 0.00 0.00 0.00 0.00
35 200 400 20470 $*{}^{*}$ 4.76 0.00 0.35 0.00 0.00 0.00 0.00 0.00
36 200 600 34504 $*{}^{*}$ 7.18 0.00 1.05 0.00 0.01 0.00 0.01 0.00
37 200 600 42860 $*{}^{*}$ 7.29 0.00 1.30 0.00 0.00 0.00 0.00 0.00
38 200 600 50984 $*{}^{*}$ 7.96 0.00 1.24 0.00 0.00 0.00 0.00 0.00
39 200 800 62625 $*{}^{*}$ 10.51 0.00 2.13 0.00 54.82 0.00 56.48 0.00
40 200 800 78387 $*{}^{*}$ 10.09 0.00 1.91 0.00 58.20 0.00 59.96 0.00
41 200 800 93978 $*{}^{*}$ 16.72 0.00 2.29 0.00 0.02 0.00 0.03 0.00
42 300 600 31000 $*{}^{*}$ 11.96 0.00 0.80 0.00 0.00 0.00 0.00 0.00
43 300 600 38216 $*{}^{*}$ 13.68 0.00 1.41 0.00 0.00 0.00 0.00 0.00
44 300 600 45310 $*{}^{*}$ 15.63 0.00 1.28 0.00 0.01 0.00 0.01 0.00
45 300 800 59600 $*{}^{*}$ 20.48 0.00 2.12 0.00 0.00 0.00 0.01 0.00
46 300 800 74500 $*{}^{*}$ 22.05 0.00 2.40 0.00 0.00 0.00 0.01 0.00
47 300 800 89300 $*{}^{*}$ 22.01 0.00 2.86 0.00 0.00 0.00 0.00 0.00
48 300 1000 96590 $*{}^{*}$ 25.59 0.00 3.10 0.00 0.02 0.00 0.02 0.00
49 300 1000 120500 $*{}^{*}$ 20.31 0.00 3.43 0.00 0.01 0.00 0.01 0.00
50 300 1000 144090 $*{}^{*}$ 23.21 0.00 3.79 0.00 0.00 0.00 0.01 0.00
Average values 9.62 0.00 1.27 0.00 7.68 0.00 8.03 0.00
%best 100 100 100 100
Table 2: Comparison of the results on ZKP Type 2 instances proposed in [35].
As far as it concerns Type 2 instances, in the literature they are solved by MEGA, and TPMT algorithms only. For this reason, Table 2 reports the results of these two algorithms as well as ClassicKS and KS. This set of instances appears very easy to solve for all the algorithms because all of them find the optimal known solution in less than 10 seconds, on average. Regarding the performance, the fastest algorithm is TPMT, which requires always less than 4 seconds, followed by MEGA with a peak of 25.59 seconds. Finally, ClassicKS and KS are slower with a peak of around 1 minute.
Figure 1: Performance comparison among HDA+, ClassicKS and KS algorithms on the CCPR instances.
Figure 2: Cumulative chart of the percentage of instances solved within a given gap at termination.
The results on the CCPR instances are reported in Table 3 in the Appendix. In this table, we only compare our approach with HDA+ method, besides ClassicKS, because the other algorithms were not tested on them. As Table 3 contains the results of 151 instances, we decided to move it into the Appendix and to represent here its results by using the cumulative charts reported in Figures 1 and 2.
The horizontal axis in Figure 1 reports the computational time in seconds, and the vertical one shows the percentage of best/optimal solutions found within that time. More precisely, a $(x,y)$ point on this plot means that within $x$ seconds of computation, the algorithm found the best/optimal solution for yof the instances in at most $x$ seconds. This implies that the faster the growth of a curve, the better the performance. It is worth noting that in our tables we compare the solutions of our algorithms with the best ones available in the literature (i.e. UB_best value). As a consequence whenever KS finds a solution which is better than UB_best, the corresponding percentage gap will be negative. Since negative percentages are not allowed in the graphic of Figure 1 then the comparison is not carried out with UB_Best, but with the best solution obtained by all the algorithms of the literature as well as our algorithms. We refer to this new value as New_UB_Best, and we will use it also for the construction of the graphic in Figure 2. The blue curve in Figure 1 is associated with the KS algorithm, the red one corresponds to ClassicKS, and the green one to the HDA+ algorithm. Figure 1 certifies the superiority of the KS with respect to the other two algorithms. Indeed, KS reaches the $\sim$68of best/optimal solutions found, while for ClassicKS and HDA+ this percentage is equal to $\sim$46and $\sim$30only, respectively. By comparing the ClassicKS and the KS curves, we observe that ClassicKS requires 262 seconds to reach its peak of $\sim$46whereas KS reaches this percentage in only 25 seconds. Moreover, after 262 seconds, KS finds the $\sim$58of best/optimal solutions which represents a relevant improvement with respect to ClassicKS. HDA+ requires only 2 seconds to reach its peak of $\sim$30but KS obtains a very close percentage (around 29%) at the same time.
Figure 2 shows the effectiveness of the three algorithms in the 151 instances considered. The horizontal axis reports the percentage gap from New_UB_Best at the end of the computation, whereas the vertical one shows the percentage of instances for which the percentage gap returned by the algorithm at termination is lower than or equal to the value reported on the horizontal axis. As for the graphic in Figure 1, the faster the growth of a curve, the better the effectiveness of the corresponding algorithm.
It is easy to see from this graphic that KS overcomes the other two algorithms because its curve is always higher than the other two ones. This means that KS obtains a lower percentage gap from the best solution, for more instances with respect to the other two algorithms.
When the horizontal value is 0%, the curves of Figure 2 show the percentage of instances for which the solution of the algorithm matches the New_UB_Best value, and this percentage is equal to $\sim$68%, $\sim$46and $\sim$30for KS, ClassicKS and HDA+, respectively. These percentages clearly coincide with the peaks reported in Figure 1 for these three algorithms. The percentage of instances for which the gap is lower than 5% significantly increases for all the algorithms. This percentage is, indeed, equal to $\sim$96%, $\sim$79and $\sim$61for KS, ClassicKS and HDA+, respectively. These results highlight the effectiveness of KS which almost always returns a solution very close to the best one and obtains a gap lower than or equal to 10% for the 98% of instances. Within the same threshold of 10%, the percentage of ClassicKS and HDA+ equals $\sim$79and $\sim$67%, respectively, which are significantly lower than the one of KS.
5 Conclusion
In this paper, we introduced a novel solution approach to the Minimum Spanning Tree with Conflict problem, leveraging the Kernel Search method [1]. This algorithm focuses on obtaining a feasible solution, ideally of high quality, by strategically selecting a small set of promising variables. This subset, known as the kernel set, is initially constructed through the solution of the LP relaxation of the original ILP. Throughout the algorithm, the kernel set is dynamically enriched by identifying new promising variables through the solution of ILP subproblems of increasing size. A key innovation in our methodology lies in the incorporation of independent sets of proper subgraphs within the KS algorithm. This novel approach draws inspiration from the observation that the edges of any feasible MSTC solution correspond to an independent set in the conflict graph.
Through comparative analysis with existing heuristics, we demonstrated the efficacy of our method in producing competitive results. It is evident that our tailored Kernel Search approach offers a promising tool for tackling the MSTC problem and potentially other conflict-based optimization challenges.
Acknowledgement
The research of F. Carrabs and M. Cerulli was partially supported by the project “SEcurity and RIghts in the CyberSpace” SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. This support is gratefully acknowledged.
Appendix
In this section, we show the detailed results of the comparison among the HDA+, ClassicKS, and KS on the CCPR instances proposed in [6]. As already noted for the other tables, whenever the B&C method finds the optimal value, we report such value with the “*” symbol in the UB_best column. Whenever no feasible solution is known in the literature, but we do find one, a “-” is reported in the UB_best column. In the gap columns, if the considered method does not find a feasible solution, we report a $100$. Instead, if the considered method does find a feasible solution, but there is a “-” in the UB_best column, we report a $-100$. Also in this case like for the previous tables, we report the results only on the instances where at least one of the methods (including ours) finds a feasible solution. As a consequence, the instances with ID 151–155, 166–170, 181, 182, 184, 185, 196–200, 211–215, 226–230 are not present in this table. The headings of Table 3 are the same as in Tables 1-2. Also in this table, we report the “Average values” and the “%best” rows at the end of the table. In both the computation of the “Average values” and the “%best” we exclude the instances in which either none of the methods in literature find a feasible solution (a “-” is reported in the UB_best column) or the KS method does not find it (the gap is 100 in its column). This happens only for four instances: 183, 191, 192, and 193.
Table 3: Comparison of the results on the CCPR instances proposed in [6], for which at least a feasible solution is found.
  Instance   HDA+   ClassicKS   KS
 ID $n$ $m$ $|C|$   UB_best   time gap  time gap  time gap
 51 25 60 18   $*{}^{*}$   0.03 0.00  0.02 0.00  0.02 0.00
 52 25 60 18   $*{}^{*}$   0.03 0.00  0.00 0.00  0.00 0.00
 53 25 60 18   $*{}^{*}$   0.03 0.00  4.15 1.98  4.47 0.00
 54 25 60 18   $*{}^{*}$   0.03 0.00  0.00 0.00  0.00 0.00
 55 25 60 18   $*{}^{*}$   0.03 0.00  0.00 0.00  0.00 0.00
 56 25 60 71   $*{}^{*}$   0.04 0.00  0.42 0.00  0.52 0.00
 57 25 60 71   $*{}^{*}$   0.04 0.00  0.62 0.26  1.19 0.26
 58 25 60 71   $*{}^{*}$   0.04 0.00  0.02 0.00  0.02 0.00
 59 25 60 71   $*{}^{*}$   0.04 0.00  0.02 0.00  0.02 0.00
 60 25 60 71   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 61 25 60 124   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 62 25 60 124   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 63 25 60 124   $*{}^{*}$   0.06 3.49  0.20 0.00  0.53 0.00
 64 25 60 124   $*{}^{*}$   0.05 0.00  0.26 1.25  0.68 0.00
 65 25 60 124   $*{}^{*}$   0.08 0.24  0.12 0.00  0.40 0.00
 66 25 90 41   $*{}^{*}$   0.03 0.00  2.80 0.00  2.84 0.00
 67 25 90 41   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 68 25 90 41   $*{}^{*}$   0.03 0.00  0.69 0.00  0.84 0.00
 69 25 90 41   $*{}^{*}$   0.03 0.00  5.08 4.38  1.89 0.00
 70 25 90 41   $*{}^{*}$   0.03 0.00  12.87 0.00  21.12 0.00
 71 25 90 161   $*{}^{*}$   0.06 0.00  0.05 0.00  0.05 0.00
 72 25 90 161   $*{}^{*}$   0.07 0.00  0.00 0.00  0.00 0.00
 73 25 90 161   $*{}^{*}$   0.06 0.00  0.02 0.00  0.02 0.00
 74 25 90 161   $*{}^{*}$   0.07 0.61  0.10 0.30  0.29 0.30
 75 25 90 161   $*{}^{*}$   0.07 0.31  0.59 1.53  1.74 0.00
 76 25 90 281   $*{}^{*}$   0.08 0.00  0.25 0.00  0.41 0.00
 77 25 90 281   $*{}^{*}$   0.10 0.00  0.91 0.00  1.22 0.00
 78 25 90 281   $*{}^{*}$   0.09 0.00  0.17 0.30  0.38 0.00
 79 25 90 281   $*{}^{*}$   0.12 2.87  0.22 0.00  0.65 0.00
 80 25 90 281   $*{}^{*}$   0.11 0.56  0.16 0.00  0.43 0.00
 81 25 120 72   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 82 25 120 72   $*{}^{*}$   0.04 0.00  0.97 0.00  7.51 0.00
 83 25 120 72   $*{}^{*}$   0.05 0.00  0.00 0.00  0.00 0.00
 84 25 120 72   $*{}^{*}$   0.04 0.00  2.56 0.00  1.18 0.00
 85 25 120 72   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 86 25 120 286   $*{}^{*}$   0.10 0.00  0.14 0.31  0.38 0.00
 87 25 120 286   $*{}^{*}$   0.11 0.00  0.51 0.00  0.53 0.00
 88 25 120 286   $*{}^{*}$   0.08 0.00  0.00 0.00  0.00 0.00
 89 25 120 286   $*{}^{*}$   0.11 0.32  0.07 0.00  0.26 0.00
 90 25 120 286   $*{}^{*}$   0.08 0.00  0.02 0.00  0.02 0.00
 91 25 120 500   $*{}^{*}$   0.17 3.65  0.71 1.82  4.49 0.61
 92 25 120 500   $*{}^{*}$   0.16 2.36  2.83 0.00  1.17 0.00
 93 25 120 500   $*{}^{*}$   0.19 4.08  1.50 0.82  0.68 0.00
 94 25 120 500   $*{}^{*}$   0.14 0.96  0.14 0.00  0.38 0.00
 95 25 120 500   $*{}^{*}$   0.16 1.25  0.19 0.62  0.58 0.62
 96 50 245 299   $*{}^{*}$   0.24 0.00  4.27 0.00  211.21 0.00
 97 50 245 299   $*{}^{*}$   0.24 0.00  0.29 0.00  1.74 0.00
 98 50 245 299   $*{}^{*}$   0.22 0.00  0.01 0.00  0.01 0.00
 99 50 245 299   $*{}^{*}$   0.26 0.00  0.35 0.16  1.57 0.00
 100 50 245 299   $*{}^{*}$   0.28 0.00  21.59 0.00  5.83 0.00
 101 50 245 1196   $*{}^{*}$   0.62 2.95  10.90 0.00  3.96 0.00
 102 50 245 1196   $*{}^{*}$   0.64 5.87  86.60 3.08  4.61 0.00
 103 50 245 1196   $*{}^{*}$   0.64 2.26  17.80 0.71  10.42 0.00
 104 50 245 1196   $*{}^{*}$   0.55 2.66  11.05 0.16  9.39 0.00
 105 50 245 1196   $*{}^{*}$   0.65 9.84  23.49 0.59  37.74 0.29
 106 50 245 2093   833   1.00 100  1102.05 1.20  456.01 -1.44
 107 50 245 2093   835   1.02 100  3685.50 100  452.01 0.00
 108 50 245 2093   840   0.91 100  2245.91 7.50  448.07 -2.26
 109 50 245 2093   836   0.97 100  1721.24 5.50  384.89 0.12
 110 50 245 2093   $*{}^{*}$   0.89 100  319.20 2.86  30.98 0.00
 111 50 367 672   $*{}^{*}$   0.57 0.00  0.49 0.35  2.62 0.00
 112 50 367 672   $*{}^{*}$   0.63 0.00  4.10 0.00  4.66 0.00
 113 50 367 672   $*{}^{*}$   0.68 0.00  0.53 0.00  2.43 0.00
 114 50 367 672   $*{}^{*}$   0.56 0.00  0.09 0.00  0.09 0.00
 115 50 367 672   $*{}^{*}$   0.64 0.36  0.44 0.18  2.59 0.00
 116 50 367 2687   $*{}^{*}$   1.52 7.35  34.96 0.33  50.68 0.00
 117 50 367 2687   $*{}^{*}$   1.44 7.80  34.21 0.81  28.74 0.00
 118 50 367 2687   $*{}^{*}$   1.50 8.18  23.60 1.53  5.18 0.34
 119 50 367 2687   $*{}^{*}$   1.54 13.72  34.00 1.74  15.30 0.00
 120 50 367 2687   $*{}^{*}$   1.57 7.00  15.72 2.33  8.92 0.00
 121 50 367 4702   726   2.26 100  3400.73 13.36  644.79 3.17
 122 50 367 4702   770   2.19 100  2418.38 16.36  870.81 0.52
 123 50 367 4702   786   2.30 100  2982.65 7.25  861.61 -0.64
 124 50 367 4702   711   2.42 100  952.21 4.92  929.33 0.28
 125 50 367 4702   764   2.16 13.61  1002.78 8.90  909.84 3.80
 126 50 490 1199   $*{}^{*}$   1.30 0.73  29.42 0.00  31.81 0.00
 127 50 490 1199   $*{}^{*}$   1.21 0.19  214.98 0.00  130.96 0.00
 128 50 490 1199   $*{}^{*}$   1.13 0.00  0.12 0.00  0.10 0.00
 129 50 490 1199   $*{}^{*}$   1.29 0.19  0.10 0.00  0.08 0.00
 130 50 490 1199   $*{}^{*}$   1.10 0.00  56.15 0.00  477.55 0.19
 131 50 490 4793   $*{}^{*}$   2.98 5.89  40.07 0.34  21.67 0.00
 132 50 490 4793   $*{}^{*}$   3.15 12.26  69.87 0.00  98.57 0.00
 133 50 490 4793   $*{}^{*}$   3.05 11.54  17.78 0.17  12.10 0.00
 134 50 490 4793   $*{}^{*}$   3.05 11.44  29.16 2.25  7.48 0.00
 135 50 490 4793   $*{}^{*}$   3.03 13.18  36.84 0.51  24.37 0.00
 136 50 490 8387   678   4.20 19.76  2445.19 16.96  449.03 -2.95
 137 50 490 8387   651   4.13 100  3606.37 15.36  964.71 0.61
 138 50 490 8387   689   4.27 100  1893.81 1.16  515.88 0.15
 139 50 490 8387   682   4.57 100  1301.15 9.09  749.00 2.05
 140 50 490 8387   674   4.47 22.85  2270.21 6.38  569.03 3.71
 141 75 555 1538   $*{}^{*}$   1.96 0.12  25.20 0.23  17.32 0.12
 142 75 555 1538   $*{}^{*}$   2.28 0.80  3.38 0.00  9.34 0.00
 143 75 555 1538   $*{}^{*}$   2.33 0.72  0.84 0.00  7.01 0.00
 144 75 555 1538   $*{}^{*}$   1.79 0.00  18.13 0.12  13.69 0.12
 145 75 555 1538   $*{}^{*}$   2.01 0.23  0.96 0.23  7.25 0.00
 146 75 555 6150   1047   5.64 18.05  2511.64 6.21  913.09 1.72
 147 75 555 6150   1069   5.57 12.91  2310.07 6.08  1038.47 -0.37
 148 75 555 6150   1040   5.50 100  3805.55 20.87  888.57 1.54
 149 75 555 6150   998   5.05 100  1086.86 1.10  543.73 -0.30
 150 75 555 6150   994   5.69 100  2127.32 18.01  1074.30 2.11
 156 75 832 3457   $*{}^{*}$   5.10 0.63  1.45 0.00  14.92 0.00
 157 75 832 3457   $*{}^{*}$   5.87 1.34  1.32 0.00  149.17 0.00
 158 75 832 3457   $*{}^{*}$   5.66 0.49  1.54 0.00  13.70 0.00
 159 75 832 3457   $*{}^{*}$   5.12 0.24  81.59 0.12  22.03 0.12
 160 75 832 3457   $*{}^{*}$   5.91 1.47  68.23 0.12  286.11 0.12
 161 75 832 13828   903   14.42 25.25  1977.61 -0.55  930.61 -0.33
 162 75 832 13828   953   13.11 18.05  3716.49 0.63  639.04 -1.57
 163 75 832 13828   892   13.52 16.03  3795.01 2.02  649.49 0.90
 164 75 832 13828   915   14.10 24.59  3343.98 1.75  1044.55 2.08
 165 75 832 13828   896   13.90 17.97  2741.60 4.24  856.37 0.89
 171 75 1110 6155   $*{}^{*}$   12.41 0.13  9.65 0.00  23.50 0.00
 172 75 1110 6155   $*{}^{*}$   10.83 0.13  50.48 0.00  51.93 0.00
 173 75 1110 6155   $*{}^{*}$   12.67 2.17  2.18 0.00  65.67 0.00
 174 75 1110 6155   $*{}^{*}$   12.30 0.64  3.04 0.13  28.39 0.00
 175 75 1110 6155   $*{}^{*}$   11.92 1.51  3.41 0.00  85.96 0.00
 176 75 1110 24620   867   27.07 19.49  2841.94 -0.35  755.04 -0.92
 177 75 1110 24620   851   29.82 23.15  3817.10 3.64  799.11 -0.12
 178 75 1110 24620   892   30.08 23.09  2969.99 6.84  1559.67 -0.90
 179 75 1110 24620   864   30.47 24.54  3356.33 2.20  895.40 0.81
 180 75 1110 24620   882   29.08 21.66  3096.31 -0.45  1572.95 -1.02
 183 75 1110 43085   1194   34.48 100  3619.07 100  3132.36 100
 186 100 990 4896   $*{}^{*}$   11.52 3.93  141.09 0.00  45.91 0.00
 187 100 990 4896   $*{}^{*}$   10.44 1.67  168.84 0.00  30.45 0.00
 188 100 990 4896   $*{}^{*}$   10.59 2.70  138.40 0.00  200.09 0.00
 189 100 990 4896   $*{}^{*}$   11.09 4.05  104.98 0.09  38.53 0.09
 190 100 990 4896   $*{}^{*}$   10.54 2.20  11.84 0.09  37.97 0.09
 191 100 990 19583   -   23.31 100  3847.28 100  1207.81 -100
 192 100 990 19583   1491   24.13 100  3832.03 100  2404.22 100
 193 100 990 19583   1510   23.93 100  3850.73 100  2060.18 100
 194 100 990 19583   1441   23.94 100  3980.25 100  1965.16 3.40
 195 100 990 19583   1560   24.58 100  3939.26 100  2594.71 5.71
 201 100 1485 11019   $*{}^{*}$   28.92 5.28  150.34 0.00  85.50 0.00
 202 100 1485 11019   $*{}^{*}$   32.30 2.65  107.84 0.00  218.25 0.00
 203 100 1485 11019   $*{}^{*}$   29.54 1.70  8.75 0.28  88.06 0.19
 204 100 1485 11019   $*{}^{*}$   30.32 1.24  115.85 1.05  142.18 0.29
 205 100 1485 11019   $*{}^{*}$   28.53 3.45  261.44 0.00  423.03 0.00
 206 100 1485 44075   1374   70.00 100  3846.93 100  3754.13 1.60
 207 100 1485 44075   1291   66.76 100  3848.22 100  3329.97 4.65
 208 100 1485 44075   1344   66.37 100  3836.87 100  3649.34 5.73
 209 100 1485 44075   1286   66.20 100  3858.32 100  3465.60 5.29
 210 100 1485 44075   1370   62.92 100  3844.58 100  3804.37 -2.26
 216 100 1980 19593   $*{}^{*}$   65.49 3.69  140.15 0.00  595.76 0.10
 217 100 1980 19593   $*{}^{*}$   70.46 4.34  95.81 0.00  118.17 0.10
 218 100 1980 19593   $*{}^{*}$   66.73 3.91  122.80 0.00  129.88 0.00
 219 100 1980 19593   $*{}^{*}$   61.12 1.76  82.52 0.00  113.50 0.00
 220 100 1980 19593   $*{}^{*}$   68.04 4.67  153.57 0.00  117.89 0.00
 221 100 1980 78369   1234   140.14 100  3893.96 100  3913.30 0.73
 222 100 1980 78369   1187   149.79 41.53  3697.46 100  3502.92 1.18
 223 100 1980 78369   1213   139.31 100  4020.74 100  3864.52 8.57
 224 100 1980 78369   1221   140.17 100  3999.16 100  2515.26 3.11
 225 100 1980 78369   $*{}^{*}$   127.94 35.42  3856.70 100  4061.45 2.25
  Average values   13.82 20.95  846.27 10.34  464.41 0.38
  %best   29.8   49.0   68.2
We notice that, for all the types of instances considered, KS returns the best average gap, as well as the highest percentage of instances with a non-positive gap with respect to the best solution value found in the literature. The 68.2% of the CCPR instances are solved with a non-positive gap, with an average gap of 0.38%, and an average computing time of less than $8$ minutes. It is definitely higher than the average computing time of the HDA+ method, which, however, returns an average gap of 20.95%. The classic KS approach is slower than KS, returning a gap of 10.34%, which is still less than the one returned by the HDA+ method.
It is clear from the comparison between the results of our KS approach and the classic KS one that using the maximum independent set problem, as well as the GRASP algorithm within the KS method, is crucial to address the MSTC problem. The increase in the computational time with respect to other existing methods is outweighed by the considerable reduction in the performance gap.